# CompilerDream: Learning a Compiler World Model for General Code Optimization

### v2

@@ Results
As shown in Table 1, CompilerDream’s policy achieves an average 1.068×1.068$\times$1.068 × code size reduction on cBench with a single trial after training, surpassing random search methods that require significantly more time—even when accounting for CompilerDream’s training time. This highlights the effectiveness of CompilerDream’s model-based RL training process in finding superior optimization passes, making it a promising component in building a general code optimizer. Moreover, when combined with guided search, CompilerDream outperforms the top method on the leaderboard while using less wall time.
4.3 General Value Prediction
We then demonstrate that the large-scale trained world model in CompilerDream can generalize to accurately simulate the optimization process of unseen input programs, thereby supporting the training of a general optimization agent. To illustrate this, we utilize our trained world model to estimate the outcome of applying an optimization sequence, a task named value prediction.
@@ 4.3 General Value Prediction
We adopt the state-of-the-art Coreset-NVP method (Liang et al. 2023), which evaluates pass sequences within a fixed candidate set (the coreset with 625 sequence prefixes) and selects the best prefix via compiler validation. Coreset-NVP employs a neural network to directly predict the normalized value of sequences, which is learned via supervised learning. Our Coreset-CompilerDream method leverages a world model trained on CodeContests, to simulate the optimization process and accumulate predicted rewards as the value. Both our method and the baseline method Coreset-NVP are tested on four datasets orthogonal to their training sets. The number of compiler validation calls is set the same across methods to ensure a fair comparison. Additional implementation details are available in Appendix B.8.
@@
Table 2: General value prediction results: The geometric mean of code size reduction achieved by the best sequences selected by different methods across 4 datasets. Results of Coreset-NVP and Coreset-Oracle is reported from Coreset-NVP’s original paper (Liang et al. 2023).
@@ Results
Table 2 shows the geometric mean reduction achieved by different methods. Despite the test datasets being distinct from the training set, CompilerDream outperforms Coreset-NVP across three datasets, highlighting its strong compiler modeling and generalization capabilities. On MiBench, CompilerDream performs nearly optimally compared to the Oracle baseline, which uses a brute-force search and serves as the problem’s upper bound. Both our method and Coreset-NVP achieve an average reduction of $1.101$ on CHStone, close to the upper bound of $1.06$.
4.4 General Code Optimization
@@
We finally aim to demonstrate that CompilerDream can learn a general optimization agent, capable of end-to-end generating effective optimization sequences for a wide variety of programs unseen during training. This scenario represents a typical daily use case where the optimization algorithm has limited time to compute the best solution for each of the diverse programs. Consequently, all methods generate only one optimization sequence per program in the test datasets unless otherwise specified.
Figure 4: Histograms of large-scale evaluations comparing CompilerDream and PPO on the FormAI dataset.
@@
Figure 5: A comparison between a ground-truth code optimization trajectory and an imagined trajectory by a learned compiler world model. The learned world model accurately captures the variations of program features and optimization metrics.
@@
Figure 6: Analysis. Evaluations of different RL algorithms: (a) Learning curves of various RL algorithms, measured by the geometric mean of code size reduction on the CodeContests validation set. A Gaussian filter ($\sigma=2.0$) is applied to enhance the visualization of trends. (b) Zero-shot generalization capabilities of different RL algorithms on various test benchmark datasets. Effect of training dataset: (c) Test performance of CompilerDream trained on CodeContests and Csmith. Bars indicate the standard deviation.
@@
We compare CompilerDream with the following baselines: a random pass sequence, an autotuning approach using random search, LLVM’s -O0 and -Oz flag, and a state-of-the-art learning-based method. The random search conducts hundreds of trials within a time budget similar to CompilerDream’s. The -Oz flag represents LLVM’s highest level of code size optimization while the -O0 flag represents no optimization. The learning-based method using Proximal Policy Optimization (PPO) (Schulman et al. 2017) and CompilerDream are trained on the large-scale CodeContests dataset and zero-shot generalized to unseen test programs. Following the state-of-the-art approach, all methods operate within a limited action space.
@@ Results
Figure 3 shows CompilerDream’s performance in reducing code size, measured by the geometric mean of IR instruction count reduction. Without specific in-domain training, CompilerDream outperforms -Oz in all but two test benchmarks in a single trial and surpasses PPO in all but the BLAS datasets. It also exceeds random search within a similar time budget on most datasets, except Linux. The minimal performance differences on the BLAS and Linux datasets suggest they are already highly optimized.
Moreover, CompilerDream’s zero-shot generalization matches or exceeds in-domain training. This advantage is especially notable in the NPB dataset, where data sparsity challenges in-domain agents, yet CompilerDream achieves an additional $3\%$ code size reduction. CompilerDream also generalizes well to new languages, evidenced in its performances on the BLAS and NPB datasets of Fortran programs. Further, on the AI-generated Objective-C dataset, CompilerDream achieves an average code size reduction of 1.027$\times$, reaching up to 2.87$\times$ in certain instances.
@@ 4.5 Analysis
We further evaluate the sample efficiency and zero-shot generalization abilities of CompilerDream based on the world model algorithm, Dreamer, against a range of model-free counterparts, including PPO (Schulman et al. 2017), DQN (Mnih et al. 2015), A2C (Mnih et al. 2016), APEX (Horgan et al. 2018), and IMPALA (Espeholt et al. 2018; Schulman et al. 2017). Figure 6a illustrates that while PPO is the strongest model-free baseline, CompilerDream outperforms it, learning an order of magnitude faster with fewer compiler interactions. As Figure 6b demonstrates, CompilerDream also excels in generalizing to unseen benchmarks, supporting our hypothesis that world model-based agents better capture the environment’s structure and enhance zero-shot generalization.
@@
To assess the impact of the CodeContests dataset on the generalization ability, we compared it with the commonly used Csmith (Yang et al. 2011) dataset, a large LLVM IR dataset generated by rules. We train CompilerDream on both datasets and the results shown in Figure 6 indicate that CompilerDream trained on CodeContests significantly outperforms the Csmith-trained version across all five test datasets. This advantage is particularly evident on the manually curated cBench (Fursin et al. 2007) and CHStone (Hara et al. 2008) datasets, demonstrating that the CodeContests-trained model can generalize more effectively to human-written programs.
@@
5 Discussion
We aim to tackle the major challenge of zero-shot generalization that arises when applying learning-based code optimization. We thus introduce the CompilerDream approach that leverages the sample efficiency and generalization capabilities of world model-based reinforcement learning, as well as the strengths of a diverse, high-quality training program dataset. Our results have shown that CompilerDream, with its powerful generalization ability, attains optimization performance that outperforms existing methods, in multiple problem scenarios under the challenging area of phase ordering. There is substantial scope for further exploration, including expansion of the training dataset, scaling up the compiler world model, optimizing multiple objectives like execution time, and enriching feature and action spaces with deeper expert knowledge or large language models.
@@ Appendix A Extended Related Work
Machine learning techniques introduced in Section 2 require crafting high-quality features that capture the important characteristics of programs, a process known as feature engineering. The most prevalent feature vectors are based on the frequencies of various types of instructions or patterns within the programs (Magni, Dubach, and O’Boyle 2014; Haj-Ali et al. 2020), designed by expert intuitions. Numerous studies have aimed to reduce the cost of feature design. Following the success of word2vec embeddings in natural language processing (Mikolov et al. 2013), methods like code2vec (Alon et al. 2019), inst2vec (Ben-Nun, Jakobovits, and Hoefler 2018), and IR2vec (VenkataKeerthy et al. 2020) represent programs as distributed vectors that capture syntactic and semantic information from the abstract syntax tree (AST) or intermediate representation (IR). The surge in deep learning has enabled feeding raw information such as AST (Alon et al. 2018), control-data flow graphs (CFG) (Brauckmann et al. 2020; Cummins et al. 2021), and code token sequences (Cummins et al. 2017) into powerful deep neural networks, capable of learning useful representations end-to-end.
@@
With the rise of large language models (LLMs), recent work has introduced LLMs into compiler optimization problems, demonstrating remarkable effectiveness in predicting optimal optimization sequences and comprehending the optimization process (Cummins et al. 2023). LLMs can directly analyze the intermediate representation of code in textual form, enabling a better understanding of the code’s intricate details. However, due to limitations on input length, it has to split programs into shorter functions for training and prediction, preventing it from predicting on code of arbitrary length. We also want to emphasize that optimizing individual functions is quite different from optimizing an entire program. Many optimization passes, such as –inline and –mergefunc, which handle function inlining and merging, are either ineffective or less critical when there’s only one function to optimize.
@@
Platforms that expose the compiler as a playground for AI experiments have significantly reduced the entry barriers to intelligent compiler research. OpenTuner (Ansel et al. 2014), and YaCoS (Filho, Rodriguez, and da Silva 2018) serve as autotuning frameworks with a range of compiler optimization techniques. Our experiments utilize CompilerGym (Cummins et al. 2022), which offers user-friendly interfaces for researchers to interact with compilers in a reinforcement learning manner. We are optimistic that the future release of our compiler world model and code optimization agents, in conjunction with these platforms, can have a democratizing effect on applying AI techniques to compiler optimizations.
@@ Appendix B Implementation Details
B.1 Compiler Environment
Our experiments are conducted on the CompilerGym platform (Cummins et al. 2022), version 0.2.5, with LLVM-10.0.0 integration.
B.2 Features and Actions
@@ 54 ArgsPhi Total arguments to Phi nodes
Table 3: Descriptions of 56-dimension Autophase features (adapted from (Cummins et al. 2022)).
We extract program features following the approach of Autophase (Haj-Ali et al. 2020). We use a 56-dimension Autophase feature vector concatenated with a 42-dimension action histogram vector as the observation. The Autophase feature contains various statistics of the IR code, with each dimension thoroughly explained in Table 3. The action histogram vector contains the counts of actions already taken by the agent within the current episode. Both vectors are normalized to ensure that their values fall within a reasonable range. Specifically, each element of the Autophase vector is divided by the total instruction count of the program, whereas the action histogram vector is normalized by the total action count per episode, set as 45.
We choose to use this feature extraction method because it is both efficient and effective. In our preliminary experiments, we find that more complex feature extraction methods, such as ProGraML (Cummins et al. 2021) and inst2vec (Ben-Nun, Jakobovits, and Hoefler 2018), significantly slow down the training and inference of CompilerDream, with only marginal improvements in performance and generalization.
We employ two distinct action spaces in our experiments to align with the baseline methods. In Sections 4.2 and Section 4.3, we utilize the full action space comprising all 124 optimization passes in LLVM. In Section 4.4 and Section 4.5, CompilerDream and the baseline methods are trained and tested on an action space derived from Autophase (Haj-Ali et al. 2020). The original version of this action space included 45 LLVM optimization passes. However, CompilerGym excludes 3 actions due to updates in the latest LLVM version. Thus, we utilize a total of 42 actions. Table 4 presents the names of the LLVM optimization passes corresponding to these 42 actions. Using this limited action space is a prevailing practice, and it has been proved to be effective in prior works(Haj-Ali et al. 2020; Cummins et al. 2022). It also makes the training easier since the model can focus on the most effective optimization passes.
@@
Table 4: A list of LLVM transformation passes selected in Autophase’s action space.
B.3 Reward Function
$r_{t+1}=\frac{C(s_{t})-C(s_{t+1})}{C(s_{0})-C(s_{\mathrm{b}})},$ (9)
where a lower $C$ indicates better performance. In line with other works focused on optimizing code size, we define $C(s)$ as the IR instruction count at compiler state $s$. The $C(s_{\mathrm{b}})$ stands for the performance of a baseline policy, specifically the built-in -Oz flag in our experiments. A total reward greater than 1 means that the optimization sequence performs better than the baseline policy.
B.4 Hyperparameters
The hyperparameters for our DreamerV3 agent implementation are outlined in Table 5. For hyperparameters not specified, we use the same value as the original DreamerV3 (Hafner et al. 2023). Most of the listed hyperparameter values are directly taken from DreamerV2 (Hafner et al. 2021) or previous works, with minimal tuning. However, the loss scales are carefully tuned, as the observation loss is relatively small compared to the reward loss in our setting, and the balance between these losses has a significant impact on CompilerDream’s performance. All experiments share the same set of hyperparameters unless otherwise specified.

### v3

@@ Results
As shown in Table 4, our RL agent trained on world model simulations achieves an average 1.068×1.068$\times$1.068 × code size reduction on cBench in a single trial, surpassing random search methods despite their longer execution times, even when accounting for CompilerDream’s training time. This demonstrates the world model’s ability to accurately simulate optimization processes and support effective agent learning. Furthermore, when combined with guided search, CompilerDream outperforms the top leaderboard method while using less wall time.
4.3. General Value Prediction
In this section, we demonstrate that CompilerDream’s large-scale trained world model accurately simulates the optimization process of unseen programs, enabling the construction of an excellent value prediction agent that surpasses state-of-the-art methods.
@@ 4.3. General Value Prediction
To ensure a fair comparison of the value prediction capabilities, we adopt the setting of the state-of-the-art Coreset-NVP method (Liang et al., 2023), which evaluates pass sequences from a fixed set of 50 distinct sequences (the core set). We train our world model to build a value prediction agent (detailed in Section 3.3) that predicts the value of applying sequences from the same core set. We follow the same search strategy as Coreset-NVP, enumerating the prefixes of roughly 3 or 4 sequences with the highest predicted value and selecting the best by compiler validation. Both methods are evaluated on four datasets distinct from their training sets, with an identical number of compiler validation calls to ensure a fair comparison. Additional details are provided in Appendix B.7.
Table 5. General value prediction results: The geometric mean of code size reduction achieved by the best sequences selected by different methods across 4 datasets.
@@ Results
Table 5 presents the geometric mean reduction achieved by different methods. Despite being tested on datasets distinct from the training set, CompilerDream surpasses Coreset-NVP on three datasets, demonstrating superior compiler simulation and generalization capabilities. On MiBench, CompilerDream performs nearly optimally compared to the Oracle baseline, which uses a brute-force search and serves as the problem’s upper bound. Both our method and Coreset-NVP achieve an average reduction of $1.101$ on CHStone, close to the upper bound of $1.06$.
4.4. General Code Optimization with RL
@@
We finally demonstrate the ability of CompilerDream’s reinforcement learning agent, which can generate optimization sequences end-to-end in a single trial for a wide variety of programs unseen during training. This scenario mirrors real-world use cases where the optimization algorithm has limited time to compute the best sequence for arbitrary programs. Therefore, all methods in this section generate only one optimization sequence per program in the test datasets unless otherwise specified.
@@
Figure 5. A comparison between a ground-truth code optimization trajectory and an imagined trajectory by a learned compiler world model. The learned world model accurately captures the variations of program features and optimization metrics.
[…]
This figure is the comparison between a ground-truth code optimization trajectory and an imagined trajectory by a learned compiler world model. Both trajectories start from the “cBench/stringsearch2” program and follow the same sequence of optimization passes (-sora, -early-cse, -simplifycfg, $\cdots$). Evaluating key metrics such as reward and intermediate representation (IR) features (BranchCount, BlockLow, TotalBlocks, $\cdots$), the world model’s predictions closely align with the real environment. After the first three passes, the mean squared error (MSE) remains low at 0.00071, 0.00088, and 0.00092, demonstrating the model’s accuracy in capturing optimization dynamics.
@@
Figure 6. Analysis. Evaluations of different RL algorithms: (a) Learning curves of various RL algorithms, measured by the geometric mean of code size reduction on the CodeContests validation set. A Gaussian filter ($\sigma=2.0$) is applied to enhance the visualization of trends. (b) Generalization capabilities of different RL algorithms on various test datasets. Effect of training dataset: (c) Test performance of CompilerDream trained on CodeContests and Csmith. Bars indicate the standard deviation.
[…]
(a)The performance of CompilerDream steadily and consistently improves throughout training, and achieves a validation code size reduction of 1.02 after about $4\times 10^{4}$ environment steps. In contrast, other methods (PPO, A2C, IMPALA, DQN, and APEX) all achieve reductions of less than 0.98 at the same stage. To reach a reduction of 1, PPO needs nearly $2\times 10^{5}$ steps–roughly six times the steps required by CompilerDream. The remaining four methods exhibit unstable performance and fail to exceed a reduction of 1 even after $10^{6}$ steps. (b) The generalization capabilities of different RL algorithms (CompilerDream, PPO, A2C, IMPALA, DQN, and APEX) on five test datasets unseen during training (cBench, CHStone, NPB, OpenCV, and TensorFlow). CompilerDream outperforms all other methods across all five datasets, demonstrating a clear advantage. PPO consistently ranks second, though with a noticeable performance gap compared to CompilerDream. The remaining four methods perform poorly, particularly on the cBench and CHStone datasets. (c) The impact of the training dataset is demonstrated by comparing CompilerDream trained on two datasets: CodeContests and Csmith. Across all five benchmarks (cBench, CHStone, NPB, OpenCV, and TensorFlow), CompilerDream trained on CodeContests consistently outperforms its counterpart trained on Csmith. Notably, the performance gap even reaches a geometric mean code size reduction difference of 0.2 on the CHStone benchmark.
Figure 7. Histograms of large-scale evaluations comparing CompilerDream and PPO on the FormAI dataset.
[…]
The performance of CompilerDream and PPO was evaluated on the large-scale FormAI test. For both methods, the code size reduction for most test cases (approximately $5\times 10^{4}$) falls within the range of $0.8\textasciitilde{}1.2$. As the reduction magnitude increases or decreases beyond this range, the number of corresponding cases gradually decreases. Comparing CompilerDream with PPO, we observe that CompilerDream optimizes more cases with higher code size reduction while producing fewer cases with worsened results. Notably, CompilerDream achieves a reduction of more than nine times for some cases, surpassing all previous studies.
@@
We train CompilerDream with its RL agent, and compare it with the following baselines: a random pass sequence, an autotuning approach using random search, LLVM’s -O0 and -Oz flag, and a state-of-the-art learning-based method (Haj-Ali et al., 2020) using Proximal Policy Optimization (PPO) (Schulman et al., 2017). The random search conducts hundreds of trials within a time budget similar to our RL agent. The -Oz flag represents LLVM’s highest level of code size optimization while the -O0 flag represents no optimization. The learning-based method using PPO and CompilerDream are trained on the large-scale CodeContests dataset and zero-shot generalized to unseen test programs. Following the state-of-the-art approach, all methods use the reduced action space (Table 1).
@@ Results
Figure 4 presents the code size reduction achieved by CompilerDream’s RL agent, measured by the geometric mean of IR instruction count reduction. Without in-domain training, CompilerDream surpasses -Oz in all but two benchmarks in a single trial and consistently outperforms PPO, except on the BLAS datasets. It also outperforms random search on most datasets within a comparable time budget, except for Linux. The minimal performance differences across various methods on BLAS and Linux suggest these datasets are already highly optimized.
Moreover, CompilerDream’s zero-shot generalization matches or surpasses in-domain training. This advantage is particularly evident in the NPB dataset, where data sparsity limits in-domain agents, yet CompilerDream achieves an additional 3% code size reduction. Its robust generalization extends to new languages, as shown by results on the Fortran-based BLAS and NPB datasets. On the AI-generated Objective-C dataset, CompilerDream achieves an average code size reduction of 1.027$\times$, reaching up to 2.87$\times$ in certain test cases.
@@ 4.5. Analysis
We further evaluate the sample efficiency and zero-shot generalization abilities of our model-based CompilerDream (with RL agent) against various model-free counterparts, including PPO (Schulman et al., 2017), DQN (Mnih et al., 2015), A2C (Mnih et al., 2016), APEX (Horgan et al., 2018), and IMPALA (Espeholt et al., 2018; Schulman et al., 2017). Figure 6a illustrates that while PPO is the strongest model-free baseline, our RL agent trained on world model simulation learns an order of magnitude faster with fewer compiler interactions. Figure 6b further highlights its superior generalization to unseen benchmarks, supporting our claim that world model-based agents better capture the dynamics of compiler optimization and enhance generalization to unseen datasets.
Comparison with model-based methods
To demonstrate the necessity and effectiveness of learning a compiler world model with deep representations of the compiler. We compare our approach with a classical model-based RL method, MBPO (Janner et al., 2019), under the RL agent setting described in Section 4.4. MBPO adopts standard MLPs for its dynamics model and does not share a latent representation space with the actor or critic. For a fair comparison, we apply reward smoothing to MBPO using the same smoothing factor as in our method. The results are shown in Table 6. While MBPO outperforms the best model-free baseline, PPO, on some tasks, it consistently underperforms CompilerDream, except on MiBench, where the performance gap across all methods remains marginal. These results suggest that model-based RL alone is insufficient, and that both a more expressive world model and a shared latent space are crucial for better optimization performance and generalization to unseen programs.
Table 6. Comparison between CompilerDream with MBPO and PPO for general code optimization.
Dataset PPO MBPO CompilerDream
BLAS 0.993 0.988 0.991
cBench 1.010 1.020 1.036
CHStone 1.071 1.066 1.094
MiBench 0.985 0.988 0.986
NPB 1.000 0.996 1.006
Linux 1.071 1.069 1.108
OpenCV 1.082 0.998 1.092
TensorFlow 1.010 0.996 1.032
Effect of reward smoothing
To evaluate the effectiveness of the reward smoothing technique described in Section 3.2, we compare the performance of the RL agent trained with the complete CompilerDream method against a variant without reward smoothing. Table 7 presents these results alongside PPO for reference. The results demonstrate that reward smoothing consistently enhances optimization performance across all datasets and enables CompilerDream to outperform PPO on MiBench, Linux, and OpenCV.
Table 7. Comparison between CompilerDream with and without reward smoothing for general code optimization.
Dataset PPO
Ours w/o
Reward Smoothing
Ours w/
Reward Smoothing
BLAS 0.993 0.987 0.991
cBench 1.010 1.021 1.036
CHStone 1.071 1.076 1.094
MiBench 1.000 0.995 1.006
NPB 1.071 1.093 1.108
Linux 0.985 0.980 0.986
OpenCV 1.082 1.080 1.092
TensorFlow 1.010 1.022 1.032
@@
To assess the impact of the CodeContests dataset on the generalization ability, we compared it with the commonly used Csmith (Yang et al., 2011) dataset, a large LLVM IR dataset generated by rules. We train CompilerDream with RL agent on both datasets and the results shown in Figure 6 indicate that CompilerDream trained on CodeContests significantly outperforms the Csmith-trained version across all five test datasets. This advantage is particularly evident in the manually curated cBench (Fursin et al., 2007) and CHStone (Hara et al., 2008) datasets, demonstrating that the CodeContests-trained model can generalize more effectively to human-written programs.
@@
5. Discussion
We aim to address the major challenge of generalization in learning-based code optimization by introducing the CompilerDream approach, which leverages the simulation and generalization capabilities of a world model. By incorporating a reward smoothing technique and a large-scale training program dataset, we enable effective world model training for compiler optimization tasks. Our method supports two types of optimization agents: a value prediction agent and an RL agent. Experimental results demonstrate that CompilerDream achieves superior optimization performance across diverse problem scenarios and program datasets, surpassing built-in compiler optimization flags and state-of-the-art methods.
Limitation
Although our method can naturally extend to optimization objectives such as execution time and object file size reduction, we focus solely on code size optimization for scalability and stability, as discussed in Section 4.1. Execution time measurements often exhibit high variance and biases due to hardware and operating system states, introducing significant noise that additional sampling cannot mitigate. Furthermore, existing open-source frameworks, such as CompilerGym, lack robust support for accurate runtime measurement, making it challenging to obtain reliable signals for effective learning. As a result, we only focus on code size reduction for this paper.
Future Works
There is substantial scope for further exploration, including expansion of the training dataset, scaling up the compiler world model, optimizing multiple objectives like execution time, and enriching feature and action spaces with deeper expert knowledge or large language models. Additionally, our approach could support more types of optimization agents, such as search agents leveraging advanced tree search algorithms (Schrittwieser et al., 2020) based on world model simulations.
6. Acknowledgments
This work was supported by the National Natural Science Foundation of China (62021002).
@@ Appendix A Extended Related Work
Machine learning techniques introduced in Section 2 require crafting high-quality features that capture the important characteristics of programs, a process known as feature engineering. The most prevalent feature vectors are based on the frequencies of various types of instructions or patterns within the programs (Magni et al., 2014; Haj-Ali et al., 2020), designed by expert intuitions. Numerous studies have aimed to reduce the cost of feature design. Following the success of word2vec embeddings in natural language processing (Mikolov et al., 2013), methods like code2vec (Alon et al., 2019), inst2vec (Ben-Nun et al., 2018), and IR2vec (VenkataKeerthy et al., 2020) represent programs as distributed vectors that capture syntactic and semantic information from the abstract syntax tree (AST) or intermediate representation (IR). The surge in deep learning has enabled feeding raw information such as AST (Alon et al., 2018), control-data flow graphs (CFG) (Brauckmann et al., 2020; Cummins et al., 2021), and code token sequences (Cummins et al., 2017) into powerful deep neural networks, capable of learning useful representations end-to-end.
@@
With the rise of large language models (LLMs), recent work has introduced LLMs into compiler optimization problems, demonstrating remarkable effectiveness in predicting optimal optimization sequences and comprehending the optimization process (Cummins et al., 2023). LLMs can directly analyze the intermediate representation of code in textual form, enabling a better understanding of the code’s intricate details. However, due to limitations on input length, it has to split programs into shorter functions for training and prediction, preventing it from predicting on code of arbitrary length. We also want to emphasize that optimizing individual functions is quite different from optimizing an entire program. Many optimization passes, such as –inline and –mergefunc, which handle function inlining and merging, are either ineffective or less critical when there’s only one function to optimize.
@@
Platforms that expose the compiler as a playground for AI experiments have significantly reduced the entry barriers to intelligent compiler research. OpenTuner (Ansel et al., 2014), and YaCoS (Filho et al., 2018) serve as autotuning frameworks with a range of compiler optimization techniques. Our experiments utilize CompilerGym (Cummins et al., 2022), which offers user-friendly interfaces for researchers to interact with compilers in a reinforcement learning manner. We are optimistic that the future release of our compiler world model and code optimization agents, in conjunction with these platforms, can have a democratizing effect on applying AI techniques to compiler optimizations.
@@ Appendix B Implementation Details
B.1. Compiler Environment
Our experiments are conducted on the CompilerGym platform (Cummins et al., 2022), version 0.2.5, with LLVM-10.0.0 integration.
B.2. Details of Features and Actions
As shown in Table 1, we adpot the 56-dimensional Autophase feature and a reduced action space from Autophase (Haj-Ali et al., 2020). A detailed list of the Autophase features and actions space is available in Table 8 and Table 9.
Table 8. Descriptions of 56-dimension Autophase features (adapted from (Cummins et al., 2022)).
@@ 54 ArgsPhi Total arguments to Phi nodes
Table 9. A list of LLVM transformation passes selected in Autophase’s action space.
@@
B.3. Hyperparameters
CompilerDream
The hyperparameters for our world model and agent implementation are outlined in Table 10. For hyperparameters not specified, we use the same value as the DreamerV3 (Hafner et al., 2023). Most of the listed hyperparameter values are directly taken from DreamerV2 (Hafner et al., 2021) or previous works, with minimal tuning. However, the loss scales are carefully tuned, as the observation loss is relatively small compared to the reward loss in our setting, and the balance between these losses has a significant impact on CompilerDream’s performance. All experiments share the same set of hyperparameters unless otherwise specified.
Model-based Baseline
For MBPO in Section 4.5, the dynamics model is a 4-layer MLP with a hidden size of 1024. We adopt the implementation from https://github.com/Xingyu-Lin/mbpo_pytorch, with most hyperparameters left unchanged. The modified hyperparameters are listed in Table 12.
Table 10. Hyperparameters in our experiments.